### v1

# RoadBench: A Vision-Language Foundation Model and Benchmark for Road Damage Understanding

Abstract
Accurate road damage detection is crucial for timely infrastructure maintenance and public safety, but existing vision-only datasets and models lack the rich contextual understanding that textual information can provide. To address this limitation, we introduce RoadBench, the first multimodal benchmark for comprehensive road damage understanding. This dataset pairs high-resolution images of road damages with detailed textual descriptions, providing a richer context for model training. We also present RoadCLIP, a novel vision-language model that builds upon CLIP by integrating domain-specific enhancements. It includes a disease-aware positional encoding that captures spatial patterns of road defects and a mechanism for injecting road-condition priors to refine the model’s understanding of road damages. We further employ a GPT-driven data generation pipeline to expand the image–text pairs in RoadBench, greatly increasing data diversity without exhaustive manual annotation. Experiments demonstrate that RoadCLIP achieves state-of-the-art performance on road damage recognition tasks, significantly outperforming existing vision-only models by 19.2%. These results highlight the advantages of integrating visual and textual information for enhanced road condition analysis, setting new benchmarks for the field and paving the way for more effective infrastructure monitoring through multimodal learning.
Figure 1: Overview of the RoadBench benchmark and the RoadCLIP framework. Left: Sample image–text pairs synthesized in diverse road scenarios, capturing damage types (e.g., longitudinal cracks, potholes), weather conditions, spatial context, and surface materials. Right: Our RoadCLIP architecture leverages a dual-encoder backbone enhanced with Disease-aware Positional Encoding and Road Disease Prior Injection to align visual and textual features in a multimodal embedding space.
1 Introduction
Road infrastructure is the backbone of economic development and social connectivity, supporting efficient transportation, commerce, and access to essential services. However, damages such as cracks, potholes, and pavement deformation significantly reduce ride quality and, if left unaddressed, pose serious safety risks, resulting in substantial economic costs. To tackle these challenges, the research community has turned to computer vision to develop automated systems for road damage detection [2, 4, 83]. The emergence of several large-scale vision-only datasets (e.g., RDD2022 [4] with over 47,000 images of road surfaces across six countries) has catalyzed the development of deep learning models for pavement distress identification and classification [6, 52, 51].
However, existing models are exclusively vision-based, relying solely on visual cues (e.g., image features), without considering textual descriptions of the damage [2, 4]. As a result, these models often struggle to capture fine-grained distinctions, such as the severity of a crack or its precise location, and may fail to generalize across diverse environmental and road conditions. Inspired by the remarkable success of multimodal approaches, like CLIP [50], which demonstrates a strong zero-shot recognition capability by learning joint image–text representations, we are motivated to explore whether multimodality could significantly improve accuracy, adaptability, and semantic understanding in real-world infrastructure monitoring scenarios. However, to our best knowledge, no publicly available dataset pairs road surface images with detailed textual descriptions, limiting the development of such multimodal solutions in this domain. Moreover, well-known large foundation models (e.g., GPT-4o [48] and DeepSeek-VL [43]) yield poor performance on road damage understanding (see Fig 6 for details), a challenge often tied to the transferability of general-purpose representations to specialized tasks [46, 87, 86].
To fill this gap, we introduce RoadBench, the first multimodal benchmark for road damage understanding, along with RoadCLIP, the first vision–language model tailored for joint image–text learning in this domain, built on this dataset. RoadBench includes 100,000 high-resolution road images, with each paired with a detailed and reliable description of the scene’s pavement condition, generated by a state-of-the-art generative language model (i.e., GPT-4o [48]). The descriptions are employed to reflect real-world conditions, including diverse environments (e.g., urban vs. rural), varied weather and lighting, and a wide range of damage types and severities. We show several examples in Figure 1. RoadCLIP comprises two key modules: a disease-aware positional encoding module and a domain-specific prior injection mechanism. The disease-aware positional encoding module injects knowledge of road geometry and common damage localization patterns into the visual branch of the multimodal backbone, making the learned representations sensitive to the road surface, where a damage occurs (e.g., wheel-path cracks and shoulder cracks), and the spatial scope of the damage, well aligned with human inspectors’ professional actions of checking both location and spread of a damage. The domain-specific prior injection equips the model with expert knowledge about typical “road diseases” (e.g., the co-occurrence of certain crack patterns or textural cues indicating material fatigue). This prior is integrated during training for image–text alignment, acting as guidance that makes multimodal representations more discriminative for our task. With the two primary innovations, RoadCLIP learns a joint embedding space where images and descriptions of road damage are tightly aligned, facilitating more accurate cross-modal understanding.
Then, we conduct extensive experiments to evaluate our dataset and model. Comparing with state-of-the-art baselines, RoadCLIP achieves superior results across multiple metrics. Notably, our approach outperforms the best purely visual model by 19.2% in detection accuracy and 20.9% in classification F1-score. It also surpasses a general vision–language baseline in image–text retrieval by a significant margin (e.g., Recall@1 improved by 14.9%). In addition, our ablation study reveals that either the positional encoding or the prior injection module plays a vital role in multimodal alignment and generalization of the proposed method.
Our primary contributions are summarized as follows:
  • We establish RoadBench, the first image-and-text benchmark for road damage understanding, with a great potential of being used for training multimodal models and evaluating methods proposed in the realm. This dataset is also the largest one (i.e., with a size of 100,000) in this field, offering diverse samples for understanding road damages.
  • We develop RoadCLIP, a new vision–language foundation model with a tailored architecture (including a novel positional encoding and prior knowledge injection) to effectively learn from road images and their descriptions. RoadCLIP achieves precise, fine-grained alignment between textual semantics and corresponding image regions (Figure 5), and demonstrates superior attention localization on road damage areas compared to GPT-4o and DeepSeek-VL (Figure 6). Moreover, it provides insights into designing domain-specific models in multimodal context.
  • We demonstrate that our approach achieves state-of-the-art performance on road damage recognition tasks, significantly outperforming existing vision-only and multimodal methods.
2 Related Work
Table 1: Comparison of RoadBench with existing single-modal road damage datasets. Our dataset is the first to provide multimodal (vision+language) annotations at scale, supporting a wide range of tasks beyond traditional detection.
Dataset Modality Samples Resolution Disease Types Task Type Availability
RDD2022 [4] Vision-only 47,420 Variable Crack, Pothole Detection Public
TD-RD [70] Vision-only 7,088 3840×2160 Crack, Pothole, Repair Detection Public
CRDDC’22 [3] Vision-only 11,720 Variable Crack, Patch, Spalling Detection Public
CNRDD [83] Vision-only 9,053 Variable Crack, Patch, Rutting Detection Public
GAPs384 [8] Vision-only 1,969 1920×1080 Crack, Pothole Detection Public
CFD [56] Vision-only 1,180 480×320 Crack Detection Public
RoadBench (Ours) Vision+Text (Multimodal) 100,000 3840×2160 Ten types in all, more details see Fig3 Multimodal Detection, Retrieval, Captioning, QA Public
2.1 Road Damage Detection Benchmarks
Existing road damage datasets and benchmarks are predominantly vision-only, focusing on detecting and localizing road surface defects from images [3, 2]. The Road Damage Dataset 2020 (RDD2020) [2] contains over 26,000 road images from multiple countries with more than 31,000 annotated damage instances, while the recent RDD2022 dataset [4] expanded this to 47,420 images across six countries, featuring over 55,000 labeled instances. Released as part of the CRDDC 2022 competition [3], these datasets have spurred the development of robust detection models using advanced architectures like YOLO series [51, 6, 78] and Transformer-based [7, 93, 90, 18, 30, 88]. Other specialized datasets include TD-RD [69], which offers aerial-view images from drone photography, GAPs384 [8] and CFD [56] for specific damage types, and CNRDD [83] featuring Chinese road conditions. However, the critical limitation across all these datasets is their unimodal nature—they contain only images and annotations without any natural language descriptions, which restricts contextual understanding, generalization capabilities [87], and semantic reasoning about damage characteristics [55, 57, 62, 34, 16, 84, 80, 35, 64, 68, 20, 67]. Detailed information on these datasets is presented in Table 1.
2.2 Vision-Language Models
There have been rapid advances in vision-language models that jointly learn from images and text [37, 29, 21, 26, 24, 14, 45, 33, 11, 54, 10, 91, 22, 63, 75, 23, 36]. Contrastive models like CLIP [50] align visual and textual representations using large-scale image–text pairs, enabling strong zero-shot performance. Unified architectures such as BLIP [31] and Florence [82] extend these capabilities to both understanding and generation through web-scale multimodal pretraining. Recent advances have further integrated powerful language models with visual processing [28, 5, 42, 38, 22, 25, 58, 65, 77, 85], including techniques like test-time adaptation [86] and prompt tuning [71, 72], as seen in models like LLaVA [40], GPT-4V [48], and Flamingo [1], which support sophisticated multimodal reasoning and contextual understanding. Evaluation typically uses benchmarks such as MS COCO Captions [39] and Flickr30k [49] for captioning and retrieval, VQAv2 [19] for visual question answering, RefCOCO [81] for referring expressions, and ScienceQA [44] for scientific reasoning. Despite this progress, a significant gap exists in domain-specific multimodal benchmarks. While similar efforts are emerging in other fields like medicine [73], none of the existing evaluation frameworks target specialized applications like road damage detection or infrastructure assessment, limiting our understanding of how these powerful models perform in critical real-world domains where both visual recognition and linguistic description are essential. .
3 Dataset Construction
In this section, we present RoadBench, the first global multimodal benchmark dataset designed for road damage detection, featuring paired high-resolution road images and corresponding detailed textual descriptions.
3.1 Data Curation
Figure 2: Overview of the RoadBench construction pipeline. Structured prompts describing road damage types and environments guide multimodal generation with GPT-4o. Human experts verify the generated image–text pairs, which are then annotated and compiled into a high-quality benchmark dataset with images, captions, and labels.
Data Generation. We firstly consult domain experts to establish a standardized vocabulary including road disease types, environmental conditions, and beyond. This expert-informed terminology guides the construction of textual prompts, which are subsequently utilized by the multimodal GPT-4o model to synthesize high-resolution (3840×2160 pixels) road surface images accompanied by corresponding textual descriptions. The generated outputs are subsequently subjected to human expert verification to ensure accuracy and relevance, a practice also found effective in other domains like medical image synthesis [61]. A mixed-method labeling approach is then applied to enhance label quality, followed by a final composition step that integrates the image, caption, and corresponding label into the RoadBench dataset.
Data Annotation. For each road damage image, we firstly annotate the location of the damage using manual labeling or using a generative approach and then generate the corresponding binary masks. These spatial annotations complement the textual descriptions, providing valuable priors for multimodal models and enabling tasks such as text-guided damage localization.
Data Validation. Civil engineering experts reviewed each image-text pair to validate the visual realism of the images and the accuracy of the corresponding textual descriptions. An iterative refinement process was also applied—flagged samples were systematically regenerated or edited to support consistent, high-quality standards across the dataset.
3.2 Data Composition
RoadBench contains 100,000 image–text pairs, each carefully generated to represent a wide range of realistic road damage scenarios. Every textual label provides precise details, including damage dimensions (e.g., a 2-meter crack), spatial positioning (e.g., center or shoulder), and environmental context (e.g., bright, wet, foggy).
Figure 3: Category-wise proportion of road defect types in RoadBench.
The dataset cover 10 major types of road surface defects — longitudinal cracks, transverse cracks, alligator cracking, potholes, patch repair, edge cracks, centerline cracks, discoloration, mixed damage patterns, and irregular/unknown defects. As shown in Figure 3, RoadBench features a reasonably balanced class distribution, with deliberate inclusion of rare yet critical defect types to simulate real-world inspection conditions and support generalized model performance. Detailed statistics of the dataset can be found in supplementary materials.
3.3 Is Synthetic Data Reliable?
Recent advances in data-driven models have highlighted the importance of large, diverse, and high-quality datasets. However, collecting and annotating real-world data is often costly, time-consuming, and privacy-sensitive, particularly in domains such as autonomous driving or road damage assessment. Consequently, synthetic data generation has emerged as a viable alternative [61, 12], enabling controlled sampling, scalable labeling, and simulation of rare or hazardous scenarios. The reliability of synthetic data has been studied across various vision tasks. For example, Richter et al. [53] demonstrated that synthetic datasets generated from video game engines (e.g., GTA V) can effectively train semantic segmentation models. Similarly, Tremblay et al. [59] showed that models trained on synthetic objects can generalize well to real-world object detection. More recently, synthetic datasets like Synscapes [66] and CARLA [15] have been widely adopted in autonomous driving research due to their realism and label precision. In this work, synthetic data serves a complementary role to real-world annotations. By leveraging a simulation pipeline augmented with text-guided prompt generation and selective augmentation, we ensure that the synthesized samples retain semantic fidelity and style diversity while remaining free of privacy concerns such as faces or license plates.
4 Methodology
Figure 4: Overall architecture of RoadCLIP. The model uses a dual-encoder CLIP-based architecture, projecting road images and damage descriptions into a shared space, trained using a symmetric contrastive loss. A Disease-aware Positional Encoding (DaPE) module adds spatial priors to the visual encoder, while a Domain-Specific Prior Injection module enriches both modalities.
4.1 Overall Architecture
RoadCLIP adopts a dual-encoder architecture building upon the CLIP framework [50], as shown in Figure 4. It consists of a vision encoder $f(\cdot)$ and a text encoder $g(\cdot)$ that project images and textual descriptions, respectively, into a shared $d$-dimensional latent space. In the vision module, we use a Transformer-based image encoder (as in CLIP, e.g. a ViT) that processes the input image as a sequence of patch embeddings. The text module is a Transformer language model encoding the input description or label. Both encoders produce a feature vector of dimension $d$ (using the Transformer’s [CLS] token or global pooled representation), which is then $\ell_{2}$-normalized and passed through a linear projection (also of dimension $d$) to produce the final image embedding $\mathbf{z}_{i}$ and text embedding $\mathbf{z}_{t}$. To adapt RoadCLIP for road-damage-aware representation learning, we incorporate two novel modules into its architecture: a Disease-aware Positional Encoding (DaPE) module integrated into the vision encoder, and a Domain-Specific Prior Injection mechanism affecting both the image and text representations.
4.2 Disease-aware Positional Encoding (DaPE)
While standard vision Transformers typically use generic positional encodings—either fixed sine-cosine patterns or learned absolute embeddings—such representations are often agnostic to the domain-specific spatial semantics needed in road damage analysis. In road scenes, spatial attributes such as crack orientation, position relative to lane markers, and edge proximity are critical for accurate classification and grounding. To this end, we propose a Disease-aware Positional Encoding (DaPE) module that encodes both absolute and geometric priors tied to road “disease” patterns.
Geometric-aware Spatial Embedding. For each image patch at normalized coordinates $(x,y)\in[0,1]^{2}$, we define a domain-specific spatial descriptor:
$\Psi_{ij}=[x,y,\cos\theta_{ij},\sin\theta_{ij}],$ (1)
where $\theta_{ij}$ represents the dominant orientation angle of any crack-like structure in patch $(i,j)$. This angle is computed using texture analysis techniques (e.g., Sobel gradients or structure tensor). The descriptor $\Psi_{ij}$ thus encodes both absolute position and directional features.
MLP-based Positional Encoding. This spatial tuple is passed through a learnable MLP $f_{\mathrm{pos}}$ to yield a $d$-dimensional positional vector:
$\mathbf{p}_{ij}^{\mathrm{(DaPE)}}=f_{\mathrm{pos}}(\Psi_{ij})\in\mathbb{R}^{d}.$ (2)
This vector is added to the patch’s visual feature embedding, either at the input stage or as a positional bias in subsequent attention layers. As a result, spatially meaningful cues—such as being near the edge or aligned along a directional crack—are encoded directly into the model’s representation.
Domain-aware Modeling. By replacing the standard positional encodings with $\mathbf{p}_{ij}^{\mathrm{(DaPE)}}$, we guide the model’s self-attention to consider spatial structures that are diagnostically meaningful. For instance, patches along a continuous horizontal crack will have similar orientation codes, encouraging them to cluster in feature space. Similarly, patches near the road edge may carry different priors than center-lane patches.
4.3 Domain-Specific Prior Injection
General-purpose vision-language models like CLIP, trained on open-world data, often lack fine-grained understanding of domain-specific semantics. To bridge this gap, RoadCLIP incorporates structured domain knowledge—specifically, road damage categories—into the training process via Domain-Specific Prior Injection, aligning visual and textual representations with semantic prototypes such as pothole, longitudinal crack, and patch repair.
Concept Embedding Initialization. We define a set of $K$ road damage types $\mathcal{C}=\{c_{1},c_{2},\dots,c_{K}\}$. Each class $c_{k}$ is associated with a joint image-text concept embedding, initialized in two ways: (1) Text-based, where a descriptive phrase (e.g., “a photo of a $c_{k}$ on a road”) is encoded by the text encoder $g(\cdot)$ to obtain $\mathbf{t}_{c_{k}}$; or (2) Learnable prototype, a trainable vector $\mathbf{v}_{k}\in\mathbb{R}^{d}$ initialized from $\mathbf{t}_{c_{k}}$. We adopt the latter, enabling the model to refine semantic priors during training while preserving interpretability.
Prior-Aligned Training Objective. For each image $I_{i}$ with damage category $y_{i}\in\mathcal{C}$, we align its projected feature $\mathbf{z}_{i}=f(I_{i})$ to its corresponding concept embedding. We define the alignment loss:
$L_{\mathrm{concept}}(I_{i})=-\log\frac{\exp(\mathrm{sim}(\mathbf{z}_{i},\mathbf{t}_{y_{i}})/\tau)}{\sum_{c\in\mathcal{C}}\exp(\mathrm{sim}(\mathbf{z}_{i},\mathbf{t}_{c})/\tau)},$ (3)
where $\mathrm{sim}(\cdot,\cdot)$ is cosine similarity and $\tau$ is a temperature parameter. This cross-entropy loss encourages image features to cluster around their concept prototypes in the embedding space.
Joint Training with Contrastive Learning. The concept alignment loss is integrated with CLIP’s standard image-text contrastive objective. Additionally, descriptive sentences for each damage type are periodically encoded to further anchor $\mathbf{v}_{k}$ in natural language space, enhancing textual generalization.
4.4 Training Objective
RoadCLIP is trained end-to-end with a composite loss that integrates contrastive alignment, concept-level supervision, and spatial robustness regularization.
Image-Text Contrastive Loss. The primary objective is a bidirectional InfoNCE loss [47] encouraging high similarity between matched image–text pairs $(I_{i},T_{i})$ and low similarity for mismatches. Let $\mathbf{z}_{i}=f(I_{i})$ and $\mathbf{z}_{i}^{+}=g(T_{i})$ denote the embeddings from the image and text encoders, respectively. The loss is:
$\mathcal{L}_{\mathrm{ITC}}=-\frac{1}{N}\sum_{i=1}^{N}\Bigl{[}\log\frac{\exp\bigl{(}\mathrm{sim}(\mathbf{z}_{i},\mathbf{z}_{i}^{+})/\tau\bigr{)}}{\sum_{j=1}^{N}\exp\bigl{(}\mathrm{sim}(\mathbf{z}_{i},\mathbf{z}_{j}^{+})/\tau\bigr{)}}\\ +\;\log\frac{\exp\bigl{(}\mathrm{sim}(\mathbf{z}_{i}^{+},\mathbf{z}_{i})/\tau\bigr{)}}{\sum_{j=1}^{N}\exp\bigl{(}\mathrm{sim}(\mathbf{z}_{j}^{+},\mathbf{z}_{i})/\tau\bigr{)}}\Bigr{]}$ (4)
Concept Prior Loss. To reinforce category-specific alignment, we apply a concept supervision loss $\mathcal{L}_{\mathrm{concept}}(I_{i})$ based on known damage type labels $y_{i}$ for all images. This loss aligns image embeddings with learnable category prototypes $\mathbf{v}_{y_{i}}$ and is weighted by $\lambda_{\mathrm{concept}}$:
$\mathcal{L}_{\mathrm{domain-align}}=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}_{\mathrm{concept}}(I_{i})$ (5)
Position Consistency Loss. To regularize DaPE, we introduce $\mathcal{L}_{\mathrm{pos-consist}}$, encouraging stability under small spatial shifts (translation/rotation) by minimizing the discrepancy between original and perturbed embeddings:
$\mathcal{L}_{\mathrm{pos-consist}}(I_{i})=\left\|f(I_{i})-f(I_{i}^{\prime})\right\|_{2}^{2}$ (6)
Total Loss. The final training loss is:
$\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{ITC}}+\lambda_{\mathrm{concept}}\mathcal{L}_{\mathrm{domain-align}}+\lambda_{\mathrm{pos}}\mathcal{L}_{\mathrm{pos-consist}}$ (7)
We optimize all model parameters (vision/text encoders, DaPE, projection layers, and concept prototypes) using Adam, with temperature $\tau$ learnable. This multi-objective training yields spatially aware and semantically aligned representations, enabling accurate road damage recognition and cross-modal retrieval.
Table 2: Comprehensive performance comparison of RoadCLIP with single-modal and multimodal state-of-the-art methods on the RoadBench dataset. Best results are highlighted in bold, and second-best results are underlined.
Single-modal Vision-only Methods Multimodal Vision-Language Models
Method Modality SLA(%) Method Modality ZS Acc.(%) Recall@1(%) Recall@5(%) Recall@10(%) SLA(%)
YOLOv10-n [NeurIPS24] [60] Vision 46.5 CLIP [ICML21] [50] Vision+Text 63.8 40.3 63.2 71.4 41.5
YOLOv10-s [NeurIPS24] [60] Vision 47.3 BLIP-2 [ICML23] [32] Vision+Text 67.9 43.7 66.5 74.0 46.3
YOLOS-ti [NeurIPS21] [17] Vision 45.3 LLaVA [NeurIPS23] [40] Vision+Text 69.1 46.0 68.9 76.2 48.0
YOLOS-s [NeurIPS21] [17] Vision 46.8 MiniGPT-4 [arXiv23] [92] Vision+Text 68.7 44.9 67.5 75.3 47.1
PP-PicoDet [arXiv21] [79] Vision 47.0 InstructBLIP [NeurIPS23] [13] Vision+Text 70.2 47.6 70.1 77.8 49.2
RT-DETR [CVPR24] [89] Vision 48.6 mPLUG-Owl [arXiv23] [76] Vision+Text 71.0 49.3 71.8 78.6 50.7
Lite-DETR [CVPR23] [30] Vision 45.3 Otter [arXiv23] [27] Vision+Text 71.4 49.7 72.2 79.1 50.9
FR-CNN [NeurIPS15] [52] Vision 39.9 GPT-4V [arXiv23] [48] Vision+Text 74.5 52.6 74.0 80.8 55.4
SSD-VGG16 [ECCV16] [41] Vision 38.7 DeepSeek-VL [arXiv24] [43] Vision+Text 72.9 50.1 72.3 79.5 53.0
TD-YOLOV10 [ICASSP25] [70] Vision 49.0 RoadCLIP (Ours) Vision+Text 78.6 58.4 78.1 84.6 61.9
5 Experiments
In this section, we present comprehensive experiments to validate the effectiveness of our proposed RoadCLIP model and the multimodal benchmark RoadBench dataset.
5.1 Experimental Setup
Datasets. We evaluate model performance primarily on our RoadBench dataset and further evaluate cross-dataset generalization on three widely used visual-only datasets: TD-RD [69], CNRDD [83], and CRDDC’22 [3].
Evaluation Metrics. Zero-Shot Classification Accuracy (ZS Acc.) measures the model’s ability to classify road damage categories without seeing labeled training data for those specific types, reflecting semantic alignment and domain adaptation capability. Image-Text Retrieval Accuracy is reported as Recall@$k$ (with $k=1,5,10$), evaluating how accurately the model retrieves the correct caption given an image (or vice versa), a standard protocol in vision-language literature [50, 32]. Additionally, we introduce Semantic Localization Accuracy (SLA) to quantify the model’s ability to correctly ground and localize defects in the image guided by textual descriptions—a task inspired by recent advances in referring expression grounding [74]. SLA is computed by measuring IoU overlap between predicted and annotated binary masks generated via text-guided attention.
Implementation Details. We firstly initialize our model with CLIP [50] pretraining weights and then fine-tune on RoadBench using a contrastive InfoNCE-style objective. Unless otherwise stated, the batch size is set to 128, learning rate to $1e^{-4}$, and training runs for 20 epochs.
5.2 Results
We compare RoadCLIP with a range of vision-only and multimodal models. As shown in Table 2, vision-only detectors such as YOLOv10 [60], PP-PicoDet [79], and RT-DETR [89] achieve moderate performance in Semantic Localization Accuracy (SLA), with TD-YOLOv10 reaching the highest score of 49.0%. However, these models lack the ability to leverage language supervision, limiting their performance in retrieval and zero-shot classification tasks. Multimodal vision-language models show stronger results across all metrics. GPT-4V [48] and DeepSeek-VL [43] achieve 74.5% and 72.9% in zero-shot accuracy (ZS Acc.), respectively, and perform competitively in image-text retrieval. In contrast, RoadCLIP achieves consistent improvements across all evaluation tasks. It obtains 78.6% in ZS Acc., outperforming GPT-4V and DeepSeek-VL by +4.1% and +5.7%, respectively. On retrieval, RoadCLIP achieves 58.4% Recall@1 and 84.6% Recall@10, establishing new state-of-the-art performance on RoadBench. In SLA, RoadCLIP surpasses GPT-4V by +6.5% and DeepSeek-VL by +8.9%, highlighting its advantage in spatial grounding. These results verify the effectiveness of our proposed Disease-aware Positional Encoding (DaPE) and domain-specific prior injection strategy. The improvements are achieved without relying on large-scale general-domain pretraining, but rather through targeted multimodal alignment tailored for the road damage domain. In summary, RoadCLIP provides a strong and reliable foundation for multimodal understanding in structured environments. We similarly provide comparative experiments with purely visual metrics, see supplementary materials for more details.
5.3 Evaluating the Effectiveness of DaPE
We evaluate the effectiveness of our Disease-aware Positional Encoding (DaPE) through an ablation study against sinusoidal, learnable absolute, relative, and no positional encoding schemes, presented in Table 3. DaPE consistently outperforms all baselines across all tasks. Specifically, DaPE improves zero-shot accuracy by +1.8%, Recall@1 by +2.5%, and SLA by a substantial +3.0% over the strongest baseline (relative encoding). These results demonstrate that incorporating domain-specific spatial priors enhances semantic alignment, generalization, and interpretability in road damage analysis. There are two additional important findings: 1) Removing positional encoding entirely leads to a significant performance drop across all metrics. The zero-shot accuracy drops to 72.1%, and SLA falls below 54%, highlighting that explicit spatial priors are critical for accurately understanding and localizing road defects. These results underscore the sensitivity of Transformer-based models to the absence of spatial signals, particularly in structured domains such as road imagery. 2) The relative position encoding consistently outperforms both sinusoidal and learnable absolute encodings. While learnable absolute encodings show moderate improvements over fixed ones, relative encodings achieve better alignment in tasks requiring spatial awareness—suggesting that modeling relationships between patches, rather than absolute positions alone, better supports semantic grounding.
Table 3: Ablation study on the effectiveness of Disease-aware Position Encoding (DaPE). Best results are highlighted in bold, and second-best results are underlined.
Configuration ZS Acc.(%) Recall@1(%) Recall@5(%) SLA(%)
RoadCLIP w/o Positional Encoding 72.1 51.0 71.5 53.7
RoadCLIP w/ Sinusoidal Absolute PE 74.2 53.8 72.8 56.1
RoadCLIP w/ Learnable Absolute PE 75.4 54.6 74.0 57.8
RoadCLIP w/ Relative PE 76.8 55.9 75.3 58.9
RoadCLIP w/ DaPE (Proposed) 78.6 58.4 78.1 61.9
Table 4: Generalization performance comparison of different multimodal models across TD-RD, CNRDD, and CRDDC’22 datasets. Best results are highlighted in bold, second-best are underlined.
Method TD-RD[70] (ZS Acc. %) CNRDD[83] (Recall@1 %) CRDDC’22[3] (SLA %)
Score $\Delta$ vs. CLIP Rank Score $\Delta$ vs. CLIP Rank Score $\Delta$ vs. CLIP Rank
CLIP [ICML21] [50] 65.7 - 6 38.2 - 6 42.5 - 6
BLIP-2 [ICML23] [32] 67.4 +1.7 5 40.6 +2.4 5 44.9 +2.4 5
LLaVA [NeurIPS23] [40] 69.1 +3.4 4 41.7 +3.5 4 47.1 +4.6 2
GPT-4V [arXiv23] [48] 71.2 +5.5 2 44.9 +6.7 2 46.8 +4.3 3
DeepSeek-VL [arXiv24] [43] 70.5 +4.8 3 43.3 +5.1 3 45.2 +2.7 4
RoadCLIP (Ours) 74.3 +8.6 1 46.9 +8.7 1 50.1 +7.6 1
5.4 Effectiveness of Multimodal Fusion
We conduct ablation experiments across single-modal and multimodal configurations to assess each modality’s contribution and the importance of domain-specific priors. The results are summarized in Table 5. RoadCLIP, which benefits from domain-specific prior injection, achieves state-of-the-art performance across all metrics. It outperforms the strongest alternative by +6.1% in ZS Acc., +8.6% in Recall@1, and +9.3% in SLA. These substantial gains underscore the importance of vision-language fusion for capturing detailed spatial semantics and improving cross-modal alignment.
Table 5: Ablation study on multimodal fusion effectiveness.
Modality Combination ZS Acc.(%) Recall@1(%) Recall@5(%) SLA(%)
Vision Only (Baseline CNN) 58.2 30.9 48.3 35.2
Vision Only (Baseline Transformer) 60.7 33.1 50.2 37.8
Text Only (Text Encoder) 48.6 - - -
Vision + Text (CLIP Original) 63.8 40.3 63.2 41.5
Vision + Text (BLIP-2 Fusion) 67.9 43.7 66.5 46.3
Vision + Text + Absolute Position Encoding 69.2 46.2 68.4 48.7
Vision + Text + Relative Position Encoding 72.5 49.8 72.9 52.6
Vision + Text + Disease Prior (RoadCLIP) 78.6 58.4 78.1 61.9
The experiments highlight three important findings: Single-modality approaches have clear limitations. Vision-only and text-only models consistently underperform, confirming that relying on a single modality is insufficient for accurately understanding road defects. Even with a Transformer backbone, the vision-only model achieves only 60.7% ZS Acc. and 33.1% Recall@1, while the text-only encoder fails to facilitate retrieval-based evaluation. These results reflect the inherent limitations and ambiguities when using either modality alone. Fusing visual and textual modalities yields significant improvements. CLIP and BLIP-2 substantially outperform unimodal baselines, demonstrating the effectiveness of multimodal alignment for road damage analysis. Specifically, BLIP-2 improves Recall@1 to 43.7% and SLA to 46.3%, validating the benefit of incorporating textual guidance into visual representation learning. Explicit positional encoding plays a crucial role. Relative positional encoding achieves notable performance gains, attaining 72.5% ZS Acc. and 49.8% Recall@1. This indicates that spatial cues—such as crack location—are essential for distinguishing fine-grained road defect categories and strengthening attention mechanisms.
5.5 Qualitative Analysis and Visualization
To assess the interpretability and alignment, we visualize RoadCLIP’s attention using Grad-CAM and cross-modal attention maps to reveal effective grounding of textual descriptions to relevant image regions.
Fine-Grained Alignment. Figure 5 illustrates the fine-grained alignment between textual semantics and corresponding image regions enabled by RoadCLIP, through the computation of similarity between textual embeddings and patch-wise image features, resulting in a cross-modal attention map that reflects semantic relevance. Unlike conventional saliency-based visualizations, this approach captures explicit alignment between high-level language semantics and visual content, thereby enabling more interpretable and controllable predictions.
Figure 5: Illustration of the text-region alignment process in RoadCLIP. RoadCLIP encodes text and image, and computes token-wise similarity between text and visual patches. This produces a cross-modal attention map that highlights semantically aligned regions in the image.
Interpretability and Robustness. Figure 6 presents comparisons of attention maps generated by RoadCLIP and baseline models. RoadCLIP produces sharper, more localized attention responses on regions indicative of road damage. This indicates that the model captures the structural cues of road damage effectively, benefiting from the integration of DaPE and domain-specific alignment objectives.
Figure 6: Explainability of RoadCLIP via heatmap++ [9]. RoadCLIP demonstrates superior attention localization on road damage regions compared to GPT-4o and DeepSeek-VL.
6 Conclusion
We introduce RoadCLIP, a novel multimodal vision-language framework tailored for fine-grained road damage analysis, along with RoadBench, the first large-scale multimodal benchmark dataset for this domain. Our approach integrates domain-specific knowledge into the CLIP backbone through two key mechanisms: Damage-aware Positional Encoding (DaPE) and Domain-Specific Prior Injection. Experiments across zero-shot classification, image-text retrieval, and semantic localization tasks demonstrate RoadCLIP’s substantial improvements over state-of-the-art models, highlighting the value of vision-language integration for infrastructure monitoring. RoadBench and RoadCLIP establish a crucial foundation for future multimodal road damage assessment research.
Acknowledgments
This manuscript was authored by Oak Ridge National Laboratory (ORNL), operated by UT-Battelle, LLC under Contract No. DE-AC05-00OR22725 with the U.S. Department of Energy. Any subjective views or opinions expressed in this paper do not necessarily represent those of the U.S. Department of Energy or the United States Government.

### v3

@@
 However, damages such as cracks, potholes, and pavement deformation significantly reduce ride quality and, if left unaddressed, pose serious safety risks, resulting in substantial economic costs.
-To tackle these challenges, the research community has turned to computer vision to develop automated systems for road damage detection [2, 4, 83].
-The emergence of several large-scale vision-only datasets (e.g., RDD2022 [4] with over 47,000 images of road surfaces across six countries) has catalyzed the development of deep learning models for pavement distress identification and classification [6, 52, 51].
-However, existing models are exclusively vision-based, relying solely on visual cues (e.g., image features), without considering textual descriptions of the damage [2, 4].
+To tackle these challenges, the research community has turned to computer vision to develop automated systems for road damage detection [arya2021rdd2020, arya2024rdd2022, zhang2022new].
+The emergence of several large-scale vision-only datasets (e.g., RDD2022 [arya2024rdd2022] with over 47,000 images of road surfaces across six countries) has catalyzed the development of deep learning models for pavement distress identification and classification [bochkovskiy2020yolov4, ren2015faster, redmon2016you].
+However, existing models are exclusively vision-based, relying solely on visual cues (e.g., image features), without considering textual descriptions of the damage [arya2021rdd2020, arya2024rdd2022].
 As a result, these models often struggle to capture fine-grained distinctions, such as the severity of a crack or its precise location, and may fail to generalize across diverse environmental and road conditions.
-Inspired by the remarkable success of multimodal approaches, like CLIP [50], which demonstrates a strong zero-shot recognition capability by learning joint image–text representations, we are motivated to explore whether multimodality could significantly improve accuracy, adaptability, and semantic understanding in real-world infrastructure monitoring scenarios.
+Inspired by the remarkable success of multimodal approaches, like CLIP [radford2021learning], which demonstrates a strong zero-shot recognition capability by learning joint image–text representations, we are motivated to explore whether multimodality could significantly improve accuracy, adaptability, and semantic understanding in real-world infrastructure monitoring scenarios.
 However, to our best knowledge, no publicly available dataset pairs road surface images with detailed textual descriptions, limiting the development of such multimodal solutions in this domain.
-Moreover, well-known large foundation models (e.g., GPT-4o [48] and DeepSeek-VL [43]) yield poor performance on road damage understanding (see Fig 6 for details), a challenge often tied to the transferability of general-purpose representations to specialized tasks [46, 87, 86].
+Moreover, well-known large foundation models (e.g., GPT-4o [openai2023gpt4] and DeepSeek-VL [lu2024deepseekvl]) yield poor performance on road damage understanding (see Fig 6 for details), a challenge often tied to the transferability of general-purpose representations to specialized tasks [NEURIPS2024_d3602fc9, zhang2025dpcore, Zhang_2025_WACV].
 To fill this gap, we introduce RoadBench, the first multimodal benchmark for road damage understanding, along with RoadCLIP, the first vision–language model tailored for joint image–text learning in this domain, built on this dataset.
-RoadBench includes 100,000 high-resolution road images, with each paired with a detailed and reliable description of the scene’s pavement condition, generated by a state-of-the-art generative language model (i.e., GPT-4o [48]).
+RoadBench includes 100,000 high-resolution road images, with each paired with a detailed and reliable description of the scene’s pavement condition, generated by a state-of-the-art generative language model (i.e., GPT-4o [openai2023gpt4]).
 The descriptions are employed to reflect real-world conditions, including diverse environments (e.g., urban vs.
@@
 Our dataset is the first to provide multimodal (vision+language) annotations at scale, supporting a wide range of tasks beyond traditional detection.
-Dataset Modality Samples Resolution Disease Types Task Type Availability
-RDD2022 [4] Vision-only 47,420 Variable Crack, Pothole Detection Public
-TD-RD [70] Vision-only 7,088 3840×2160 Crack, Pothole, Repair Detection Public
-CRDDC’22 [3] Vision-only 11,720 Variable Crack, Patch, Spalling Detection Public
-CNRDD [83] Vision-only 9,053 Variable Crack, Patch, Rutting Detection Public
-GAPs384 [8] Vision-only 1,969 1920×1080 Crack, Pothole Detection Public
-CFD [56] Vision-only 1,180 480×320 Crack Detection Public
-RoadBench (Ours) Vision+Text (Multimodal) 100,000 3840×2160 Ten types in all, more details see Fig3 Multimodal Detection, Retrieval, Captioning, QA Public
+[…]
+RDD2022 [arya2024rdd2022] Vision-only 47,420 Variable Crack, Pothole Detection Public
+TD-RD [xiao2025tdrdtopdownbenchmarkrealtime] Vision-only 7,088 3840×2160 Crack, Pothole, Repair Detection Public
+CRDDC’22 [arya2022crowdsensing] Vision-only 11,720 Variable Crack, Patch, Spalling Detection Public
+CNRDD [zhang2022new] Vision-only 9,053 Variable Crack, Patch, Rutting Detection Public
+GAPs384 [cha2024deep] Vision-only 1,969 1920×1080 Crack, Pothole Detection Public
+CFD [shi2016automatic] Vision-only 1,180 480×320 Crack Detection Public
+[…]
 2.1 Road Damage Detection Benchmarks
-Existing road damage datasets and benchmarks are predominantly vision-only, focusing on detecting and localizing road surface defects from images [3, 2].
-The Road Damage Dataset 2020 (RDD2020) [2] contains over 26,000 road images from multiple countries with more than 31,000 annotated damage instances, while the recent RDD2022 dataset [4] expanded this to 47,420 images across six countries, featuring over 55,000 labeled instances.
-Released as part of the CRDDC 2022 competition [3], these datasets have spurred the development of robust detection models using advanced architectures like YOLO series [51, 6, 78] and Transformer-based [7, 93, 90, 18, 30, 88].
-Other specialized datasets include TD-RD [69], which offers aerial-view images from drone photography, GAPs384 [8] and CFD [56] for specific damage types, and CNRDD [83] featuring Chinese road conditions.
-However, the critical limitation across all these datasets is their unimodal nature—they contain only images and annotations without any natural language descriptions, which restricts contextual understanding, generalization capabilities [87], and semantic reasoning about damage characteristics [55, 57, 62, 34, 16, 84, 80, 35, 64, 68, 20, 67].
+Existing road damage datasets and benchmarks are predominantly vision-only, focusing on detecting and localizing road surface defects from images [arya2022crowdsensing, arya2021rdd2020].
+The Road Damage Dataset 2020 (RDD2020) [arya2021rdd2020] contains over 26,000 road images from multiple countries with more than 31,000 annotated damage instances, while the recent RDD2022 dataset [arya2024rdd2022] expanded this to 47,420 images across six countries, featuring over 55,000 labeled instances.
+Released as part of the CRDDC 2022 competition [arya2022crowdsensing], these datasets have spurred the development of robust detection models using advanced architectures like YOLO series [redmon2016you, bochkovskiy2020yolov4, yu2021pp] and Transformer-based [carion2020end, zhu2021deformable, zhao2024detrs, fang2021you, li2023lite, zhang2021vit].
+Other specialized datasets include TD-RD [xiao2025tdrd], which offers aerial-view images from drone photography, GAPs384 [cha2024deep] and CFD [shi2016automatic] for specific damage types, and CNRDD [zhang2022new] featuring Chinese road conditions.
+However, the critical limitation across all these datasets is their unimodal nature—they contain only images and annotations without any natural language descriptions, which restricts contextual understanding, generalization capabilities [zhang2025dpcore], and semantic reasoning about damage characteristics [sharma2018conceptual, subramanian2020medicat, wang2024ensemble, li2024yolox, dugalam2024development, zhang2024real, yu2024road, 9904187, wang2024drivedreamer, wu2024seesr, he2024can, wu2024one].
 Detailed information on these datasets is presented in Table 1.
 2.2 Vision-Language Models
-There have been rapid advances in vision-language models that jointly learn from images and text [37, 29, 21, 26, 24, 14, 45, 33, 11, 54, 10, 91, 22, 63, 75, 23, 36].
-Contrastive models like CLIP [50] align visual and textual representations using large-scale image–text pairs, enabling strong zero-shot performance.
-Unified architectures such as BLIP [31] and Florence [82] extend these capabilities to both understanding and generation through web-scale multimodal pretraining.
-Recent advances have further integrated powerful language models with visual processing [28, 5, 42, 38, 22, 25, 58, 65, 77, 85], including techniques like test-time adaptation [86] and prompt tuning [71, 72], as seen in models like LLaVA [40], GPT-4V [48], and Flamingo [1], which support sophisticated multimodal reasoning and contextual understanding.
-Evaluation typically uses benchmarks such as MS COCO Captions [39] and Flickr30k [49] for captioning and retrieval, VQAv2 [19] for visual question answering, RefCOCO [81] for referring expressions, and ScienceQA [44] for scientific reasoning.
+There have been rapid advances in vision-language models that jointly learn from images and text [li2025benchmark, li2022elevater, hong2020more, kuckreja2024geochat, jain2024semantic, dong2024changeclip, marino2019ok, li2024survey, chen2023benchmarking, roberts2024image2struct, chen2024viseval, zheng2022vlmbench, hong2024cogagent, wang2024cogvlm, yao2024tcp, huang2024adapting, li2024cycleyoloefficientrobustframework].
+Contrastive models like CLIP [radford2021learning] align visual and textual representations using large-scale image–text pairs, enabling strong zero-shot performance.
+Unified architectures such as BLIP [li2022blip] and Florence [yuan2021florence] extend these capabilities to both understanding and generation through web-scale multimodal pretraining.
+Recent advances have further integrated powerful language models with visual processing [li2024seed, bitton2023visit, liu2024ii, lin2024vila, hong2024cogagent, kayser2021vil, tankala2024wikido, wang2024journeybench, yin2023lamm, zhang2024unveiling], including techniques like test-time adaptation [Zhang_2025_WACV] and prompt tuning [xiao2025visualinstanceawareprompttuning, xiao2025visualvariationalautoencoderprompt], as seen in models like LLaVA [liu2023llava], GPT-4V [openai2023gpt4], and Flamingo [alayrac2022flamingo], which support sophisticated multimodal reasoning and contextual understanding.
+Evaluation typically uses benchmarks such as MS COCO Captions [lin2014microsoft] and Flickr30k [plummer2015flickr30k] for captioning and retrieval, VQAv2 [goyal2017making] for visual question answering, RefCOCO [yu2016modeling] for referring expressions, and ScienceQA [lu2022learn] for scientific reasoning.
 Despite this progress, a significant gap exists in domain-specific multimodal benchmarks.
-While similar efforts are emerging in other fields like medicine [73], none of the existing evaluation frameworks target specialized applications like road damage detection or infrastructure assessment, limiting our understanding of how these powerful models perform in critical real-world domains where both visual recognition and linguistic description are essential.
+While similar efforts are emerging in other fields like medicine [xiao2025medicalimages], none of the existing evaluation frameworks target specialized applications like road damage detection or infrastructure assessment, limiting our understanding of how these powerful models perform in critical real-world domains where both visual recognition and linguistic description are essential.
 .
@@
 This expert-informed terminology guides the construction of textual prompts, which are subsequently utilized by the multimodal GPT-4o model to synthesize high-resolution (3840×2160 pixels) road surface images accompanied by corresponding textual descriptions.
-The generated outputs are subsequently subjected to human expert verification to ensure accuracy and relevance, a practice also found effective in other domains like medical image synthesis [61].
+The generated outputs are subsequently subjected to human expert verification to ensure accuracy and relevance, a practice also found effective in other domains like medical image synthesis [wang2025doctor].
 A mixed-method labeling approach is then applied to enhance label quality, followed by a final composition step that integrates the image, caption, and corresponding label into the RoadBench dataset.
@@
 However, collecting and annotating real-world data is often costly, time-consuming, and privacy-sensitive, particularly in domains such as autonomous driving or road damage assessment.
-Consequently, synthetic data generation has emerged as a viable alternative [61, 12], enabling controlled sampling, scalable labeling, and simulation of rare or hazardous scenarios.
+Consequently, synthetic data generation has emerged as a viable alternative [wang2025doctor, chung2025soksyntheticimagesreplace], enabling controlled sampling, scalable labeling, and simulation of rare or hazardous scenarios.
 The reliability of synthetic data has been studied across various vision tasks.
 For example, Richter et al.
-[53] demonstrated that synthetic datasets generated from video game engines (e.g., GTA V) can effectively train semantic segmentation models.
+[richter2016playing] demonstrated that synthetic datasets generated from video game engines (e.g., GTA V) can effectively train semantic segmentation models.
 Similarly, Tremblay et al.
-[59] showed that models trained on synthetic objects can generalize well to real-world object detection.
-More recently, synthetic datasets like Synscapes [66] and CARLA [15] have been widely adopted in autonomous driving research due to their realism and label precision.
+[tremblay2018training] showed that models trained on synthetic objects can generalize well to real-world object detection.
+More recently, synthetic datasets like Synscapes [wrenninge2018synscapes] and CARLA [dosovitskiy2017carla] have been widely adopted in autonomous driving research due to their realism and label precision.
 In this work, synthetic data serves a complementary role to real-world annotations.
@@
 4.1 Overall Architecture
-RoadCLIP adopts a dual-encoder architecture building upon the CLIP framework [50], as shown in Figure 4.
+RoadCLIP adopts a dual-encoder architecture building upon the CLIP framework [radford2021learning], as shown in Figure 4.
 It consists of a vision encoder $f(\cdot)$ and a text encoder $g(\cdot)$ that project images and textual descriptions, respectively, into a shared $d$-dimensional latent space.
@@
 Image-Text Contrastive Loss.
-The primary objective is a bidirectional InfoNCE loss [47] encouraging high similarity between matched image–text pairs $(I_{i},T_{i})$ and low similarity for mismatches.
+The primary objective is a bidirectional InfoNCE loss [oord2018representation] encouraging high similarity between matched image–text pairs $(I_{i},T_{i})$ and low similarity for mismatches.
 Let $\mathbf{z}_{i}=f(I_{i})$ and $\mathbf{z}_{i}^{+}=g(T_{i})$ denote the embeddings from the image and text encoders, respectively.
@@
 Best results are highlighted in bold, and second-best results are underlined.
-Single-modal Vision-only Methods Multimodal Vision-Language Models
-Method Modality SLA(%) Method Modality ZS Acc.(%) Recall@1(%) Recall@5(%) Recall@10(%) SLA(%)
-YOLOv10-n [NeurIPS24] [60] Vision 46.5 CLIP [ICML21] [50] Vision+Text 63.8 40.3 63.2 71.4 41.5
-YOLOv10-s [NeurIPS24] [60] Vision 47.3 BLIP-2 [ICML23] [32] Vision+Text 67.9 43.7 66.5 74.0 46.3
-YOLOS-ti [NeurIPS21] [17] Vision 45.3 LLaVA [NeurIPS23] [40] Vision+Text 69.1 46.0 68.9 76.2 48.0
-YOLOS-s [NeurIPS21] [17] Vision 46.8 MiniGPT-4 [arXiv23] [92] Vision+Text 68.7 44.9 67.5 75.3 47.1
-PP-PicoDet [arXiv21] [79] Vision 47.0 InstructBLIP [NeurIPS23] [13] Vision+Text 70.2 47.6 70.1 77.8 49.2
-RT-DETR [CVPR24] [89] Vision 48.6 mPLUG-Owl [arXiv23] [76] Vision+Text 71.0 49.3 71.8 78.6 50.7
-Lite-DETR [CVPR23] [30] Vision 45.3 Otter [arXiv23] [27] Vision+Text 71.4 49.7 72.2 79.1 50.9
-FR-CNN [NeurIPS15] [52] Vision 39.9 GPT-4V [arXiv23] [48] Vision+Text 74.5 52.6 74.0 80.8 55.4
-SSD-VGG16 [ECCV16] [41] Vision 38.7 DeepSeek-VL [arXiv24] [43] Vision+Text 72.9 50.1 72.3 79.5 53.0
-TD-YOLOV10 [ICASSP25] [70] Vision 49.0 RoadCLIP (Ours) Vision+Text 78.6 58.4 78.1 84.6 61.9
+[…]
+[…]
+[…]
+YOLOv10-s [NeurIPS24] [wang2024yolov10] Vision 47.3 BLIP-2 [ICML23] [li2023blip] Vision+Text 67.9 43.7 66.5 74.0 46.3
+YOLOS-ti [NeurIPS21] [Fang2021YouOL] Vision 45.3 LLaVA [NeurIPS23] [liu2023llava] Vision+Text 69.1 46.0 68.9 76.2 48.0
+YOLOS-s [NeurIPS21] [Fang2021YouOL] Vision 46.8 MiniGPT-4 [arXiv23] [zhu2023minigpt] Vision+Text 68.7 44.9 67.5 75.3 47.1
+PP-PicoDet [arXiv21] [yu2021pppicodetbetterrealtimeobject] Vision 47.0 InstructBLIP [NeurIPS23] [dai2023instructblip] Vision+Text 70.2 47.6 70.1 77.8 49.2
+RT-DETR [CVPR24] [Zhao_2024_CVPR] Vision 48.6 mPLUG-Owl [arXiv23] [ye2023mplug] Vision+Text 71.0 49.3 71.8 78.6 50.7
+Lite-DETR [CVPR23] [li2023lite] Vision 45.3 Otter [arXiv23] [li2023ottermultimodalmodelincontext] Vision+Text 71.4 49.7 72.2 79.1 50.9
+FR-CNN [NeurIPS15] [ren2015faster] Vision 39.9 GPT-4V [arXiv23] [openai2023gpt4] Vision+Text 74.5 52.6 74.0 80.8 55.4
+SSD-VGG16 [ECCV16] [liu2016ssd] Vision 38.7 DeepSeek-VL [arXiv24] [lu2024deepseekvl] Vision+Text 72.9 50.1 72.3 79.5 53.0
+TD-YOLOV10 [ICASSP25] [xiao2025tdrdtopdownbenchmarkrealtime] Vision 49.0 RoadCLIP (Ours) Vision+Text 78.6 58.4 78.1 84.6 61.9
 5 Experiments
@@
 Datasets.
-We evaluate model performance primarily on our RoadBench dataset and further evaluate cross-dataset generalization on three widely used visual-only datasets: TD-RD [69], CNRDD [83], and CRDDC’22 [3].
+We evaluate model performance primarily on our RoadBench dataset and further evaluate cross-dataset generalization on three widely used visual-only datasets: TD-RD [xiao2025tdrd], CNRDD [zhang2022new], and CRDDC’22 [arya2022crowdsensing].
 Evaluation Metrics.
 Zero-Shot Classification Accuracy (ZS Acc.) measures the model’s ability to classify road damage categories without seeing labeled training data for those specific types, reflecting semantic alignment and domain adaptation capability.
-Image-Text Retrieval Accuracy is reported as Recall@$k$ (with $k=1,5,10$), evaluating how accurately the model retrieves the correct caption given an image (or vice versa), a standard protocol in vision-language literature [50, 32].
-Additionally, we introduce Semantic Localization Accuracy (SLA) to quantify the model’s ability to correctly ground and localize defects in the image guided by textual descriptions—a task inspired by recent advances in referring expression grounding [74].
+Image-Text Retrieval Accuracy is reported as Recall@$k$ (with $k=1,5,10$), evaluating how accurately the model retrieves the correct caption given an image (or vice versa), a standard protocol in vision-language literature [radford2021learning, li2023blip].
+Additionally, we introduce Semantic Localization Accuracy (SLA) to quantify the model’s ability to correctly ground and localize defects in the image guided by textual descriptions—a task inspired by recent advances in referring expression grounding [yang2023semantics].
 SLA is computed by measuring IoU overlap between predicted and annotated binary masks generated via text-guided attention.
 Implementation Details.
-We firstly initialize our model with CLIP [50] pretraining weights and then fine-tune on RoadBench using a contrastive InfoNCE-style objective.
+We firstly initialize our model with CLIP [radford2021learning] pretraining weights and then fine-tune on RoadBench using a contrastive InfoNCE-style objective.
 Unless otherwise stated, the batch size is set to 128, learning rate to $1e^{-4}$, and training runs for 20 epochs.
@@
 We compare RoadCLIP with a range of vision-only and multimodal models.
-As shown in Table 2, vision-only detectors such as YOLOv10 [60], PP-PicoDet [79], and RT-DETR [89] achieve moderate performance in Semantic Localization Accuracy (SLA), with TD-YOLOv10 reaching the highest score of 49.0%.
+As shown in Table 4.4, vision-only detectors such as YOLOv10 [wang2024yolov10], PP-PicoDet [yu2021pppicodetbetterrealtimeobject], and RT-DETR [Zhao_2024_CVPR] achieve moderate performance in Semantic Localization Accuracy (SLA), with TD-YOLOv10 reaching the highest score of 49.0%.
 However, these models lack the ability to leverage language supervision, limiting their performance in retrieval and zero-shot classification tasks.
 Multimodal vision-language models show stronger results across all metrics.
-GPT-4V [48] and DeepSeek-VL [43] achieve 74.5% and 72.9% in zero-shot accuracy (ZS Acc.), respectively, and perform competitively in image-text retrieval.
+GPT-4V [openai2023gpt4] and DeepSeek-VL [lu2024deepseekvl] achieve 74.5% and 72.9% in zero-shot accuracy (ZS Acc.), respectively, and perform competitively in image-text retrieval.
 In contrast, RoadCLIP achieves consistent improvements across all evaluation tasks.
@@
 Table 3: Ablation study on the effectiveness of Disease-aware Position Encoding (DaPE). Best results are highlighted in bold, and second-best results are underlined.
-Configuration ZS Acc.(%) Recall@1(%) Recall@5(%) SLA(%)
+[…]
 RoadCLIP w/o Positional Encoding 72.1 51.0 71.5 53.7
@@
 Table 4: Generalization performance comparison of different multimodal models across TD-RD, CNRDD, and CRDDC’22 datasets. Best results are highlighted in bold, second-best are underlined.
-Method TD-RD[70] (ZS Acc. %) CNRDD[83] (Recall@1 %) CRDDC’22[3] (SLA %)
-Score $\Delta$ vs. CLIP Rank Score $\Delta$ vs. CLIP Rank Score $\Delta$ vs. CLIP Rank
-CLIP [ICML21] [50] 65.7 - 6 38.2 - 6 42.5 - 6
-BLIP-2 [ICML23] [32] 67.4 +1.7 5 40.6 +2.4 5 44.9 +2.4 5
-LLaVA [NeurIPS23] [40] 69.1 +3.4 4 41.7 +3.5 4 47.1 +4.6 2
-GPT-4V [arXiv23] [48] 71.2 +5.5 2 44.9 +6.7 2 46.8 +4.3 3
-DeepSeek-VL [arXiv24] [43] 70.5 +4.8 3 43.3 +5.1 3 45.2 +2.7 4
+[…]
+[…]
+[…]
+BLIP-2 [ICML23] [li2023blip] 67.4 +1.7 5 40.6 +2.4 5 44.9 +2.4 5
+LLaVA [NeurIPS23] [liu2023llava] 69.1 +3.4 4 41.7 +3.5 4 47.1 +4.6 2
+GPT-4V [arXiv23] [openai2023gpt4] 71.2 +5.5 2 44.9 +6.7 2 46.8 +4.3 3
+DeepSeek-VL [arXiv24] [lu2024deepseekvl] 70.5 +4.8 3 43.3 +5.1 3 45.2 +2.7 4
 RoadCLIP (Ours) 74.3 +8.6 1 46.9 +8.7 1 50.1 +7.6 1
-5.4 Effectiveness of Multimodal Fusion
-We conduct ablation experiments across single-modal and multimodal configurations to assess each modality’s contribution and the importance of domain-specific priors.
-The results are summarized in Table 5.
-RoadCLIP, which benefits from domain-specific prior injection, achieves state-of-the-art performance across all metrics.
-It outperforms the strongest alternative by +6.1% in ZS Acc., +8.6% in Recall@1, and +9.3% in SLA.
-These substantial gains underscore the importance of vision-language fusion for capturing detailed spatial semantics and improving cross-modal alignment.
-Table 5: Ablation study on multimodal fusion effectiveness.
-Modality Combination ZS Acc.(%) Recall@1(%) Recall@5(%) SLA(%)
-Vision Only (Baseline CNN) 58.2 30.9 48.3 35.2
-Vision Only (Baseline Transformer) 60.7 33.1 50.2 37.8
-Text Only (Text Encoder) 48.6 - - -
-Vision + Text (CLIP Original) 63.8 40.3 63.2 41.5
-Vision + Text (BLIP-2 Fusion) 67.9 43.7 66.5 46.3
-Vision + Text + Absolute Position Encoding 69.2 46.2 68.4 48.7
-Vision + Text + Relative Position Encoding 72.5 49.8 72.9 52.6
-Vision + Text + Disease Prior (RoadCLIP) 78.6 58.4 78.1 61.9
-The experiments highlight three important findings: Single-modality approaches have clear limitations.
-Vision-only and text-only models consistently underperform, confirming that relying on a single modality is insufficient for accurately understanding road defects.
-Even with a Transformer backbone, the vision-only model achieves only 60.7% ZS Acc.
-and 33.1% Recall@1, while the text-only encoder fails to facilitate retrieval-based evaluation.
-These results reflect the inherent limitations and ambiguities when using either modality alone.
-Fusing visual and textual modalities yields significant improvements.
-CLIP and BLIP-2 substantially outperform unimodal baselines, demonstrating the effectiveness of multimodal alignment for road damage analysis.
-Specifically, BLIP-2 improves Recall@1 to 43.7% and SLA to 46.3%, validating the benefit of incorporating textual guidance into visual representation learning.
-Explicit positional encoding plays a crucial role.
-Relative positional encoding achieves notable performance gains, attaining 72.5% ZS Acc.
-and 49.8% Recall@1.
-This indicates that spatial cues—such as crack location—are essential for distinguishing fine-grained road defect categories and strengthening attention mechanisms.
 5.4 Qualitative Analysis and Visualization
@@
 This indicates that the model captures the structural cues of road damage effectively, benefiting from the integration of DaPE and domain-specific alignment objectives.
-Figure 6: Explainability of RoadCLIP via heatmap++ [9]. RoadCLIP demonstrates superior attention localization on road damage regions compared to GPT-4o and DeepSeek-VL.
+Figure 6: Explainability of RoadCLIP via heatmap++ [chefer2022attentive]. RoadCLIP demonstrates superior attention localization on road damage regions compared to GPT-4o and DeepSeek-VL.
 6 Conclusion
@@
 RoadBench and RoadCLIP establish a crucial foundation for future multimodal road damage assessment research.
-Acknowledgments
-This manuscript was authored by Oak Ridge National Laboratory (ORNL), operated by UT-Battelle, LLC under Contract No.
+Acknowledgment
+This manuscript was co-authored by Oak Ridge National Laboratory (ORNL), operated by UT-Battelle, LLC under Contract No.
 DE-AC05-00OR22725 with the U.S.